ELECTRONIC SPECTROSCOPY OF JET-COOLED RHODIUM MONONITRIDE AND RHODIUM MONO-CARBIDE IN THE VISIBLE REGION

\underline{S. FOUGERE}, W. J. BALFOUR, R. F. HEUFF, and C. X. W. QIAN, Department of Chemistry, University of Victoria, Victoria, B.C. Canada V8W 3P6.

We report the gas phase spectroscopic studies of the rhodium mononitride (RhN) and rhodium monocarbide (RhC) radicals. RhN/RhC molecules were produced in a laser vaporization apparatus utilizing rhodium metal and ammonia/methane. Laser induced fluorescence spectra were recorded in the 390-750 nm region. It is believed that this is the first observation of the RhN radical in the laboratory. Transitions with Λ = 0 and 1 have been noted. Rotational and vibrational analyses have been performed and assignments made based on these and on ^{14}N/^{15}N isotopic shifts, excited state lifetimes and dispersed fluorescence (DF) spectra. The Rh^{14}N DF spectra reveal the presence of several low-lying electronic states and long progressions have provided information on the vibrational spacing. A ...δ^{4}σ^{2} electronic configuration is proposed for the ground state of RhN which agrees with the observation that all the jet-cooled spectra have ^{1}Σ as their lower state. Possible electronic configurations for other low-lying electronic states as well as excited states will be discussed. The RhC bands observed by Lagerqvist and Scullman^{a} have been re-examined. Many bands not reported by them have also been observed. The new results suggest that the B-X spectra need to be reinterpreted. The DF spectra from both B and C states show long ground state progressions(v=0 to v=10). Emission to the A^{2}Π state has also been observed.

aA. Lagerqvist and R. Scullman, Ark. Fys., 32, 479 (1965).